Triacylglycerol A contains 53% oleic acid, 29% linoleic acid, 9% palmitic acid, 4% stearic acid, 1% myristic acid and lesser amounts of other fatty acids. Triacylglycerol B contains 39% oleic acid, 29% palmitic acid, 24% stearic acid, 2% myristic acid, 2% linoleic acid, and lesser amounts of other fatty acids. Which of the two samples is expected to be solid at room temperature? Why pls? Thanks
Answer #1

Here, oleic acid and linoleic acid are the unsaturated fatty acids, hence they are liquids.

Whereas palmitic acid, stearic acid, and myristic acid are the saturated fatty acids, hence they are solids.

The solid composition is high in the case of Triacylglycerol A = 9+4+1 = 13%

The solid composition is high in the case of Triacylglycerol B = 29+24+2 = 55%

Hence, Triacylglycerol B is expected to be solid at room temperature.